OpenStudy (anonymous):

(1/2 + 1/4) divided by (1/2 + 1/3)

OpenStudy (anonymous):

answer is 1/2

OpenStudy (anonymous):

i got 9/10 for the answer. how did you get 1/2?

OpenStudy (anonymous):

i cancel directly 1/2 and 1/2, then i have the remaining is 1/4 divided by 1/3 then get the reciprocal of 1/3 its 3, sorry the answer is 3/4.

OpenStudy (anonymous):

what in the world. you have to add, then divide: \[(\frac{1}{2}+\frac{1}{4})\div (\frac{1}{2} + \frac{1}{3})=\frac{3}{4}\div \frac{5}{6}\] \[=\frac{3}{4}\times \frac{6}{5}=\frac{18}{20}=\frac{9}{10}\]